THE ν_{2} BENDING VIBRATIONAL STRUCTURE OF THE X̃ ^{2}Σ^{+} STATE OF MgNC

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated MgNC in supersonic free jet expansions, and observed the laser induced fluorescence ( LIF ) of the Ã 2Π – X̃ ^{2}Σ^{+} transition. We measured the LIF dispersed fluorescence spectra from the single vibronic levels ( SVL ) on the Mg-N-C bending ( ν_{2} ) mode in the upper Ã^{2}Π state. On the basis of the vibronic structures in the SVL dispersed spectra, we discuss the vibrational structure of the ground X̃ ^{2}Σ^{+} state of MgNC and the potential surface of the isomerization reaction, MgNC ↔ MgCN.